# Edge of Stochastic Stability: Revisiting the Edge of Stability for SGD

### v3

@@ Appendix B Stability for Mini-Batch SGD
Note that the Taylor expansion as in Equation 10 ensures that
$\mathbb{E}[f_{B}(x_{t+1})]\leq\mathbb{E}[f_{B}(x_{t})]\ =\ \mathbb{E}\left[-\eta\nabla f_{B}(x_{t})^{\top}\nabla f_{B}(x_{t})\,+\,\frac{\eta^{2}}{2}\nabla f_{B}(x_{t})^{\top}\cdot\nabla^{2}f_{B}(x_{t})\nabla f_{B}(x_{t})\right]$
Thus, by enforcing $\mathbb{E}[f_{B}(x_{t+1})]\leq\mathbb{E}[f_{B}(x_{t})]$ is equivalent to enforcing
$2\ \geq\ \eta\cdot\textit{Batch Sharpness}.$
Taking this expectation over the ways in which we can sample the batches we conclude the proof. ∎
Appendix C Comparison with Lee & Jang (2023)
After we wrote the first version of this article, we realized that Lee & Jang (2023) already dealt with this problem. We draw here a comparison and show that they come to a different conclusion because they deal with a different notion of stability. While our batch sharpness and their Interaction-Aware Sharpness (IAS) are similar in the experiments–and about the same in case of bigger batch sizes, we discuss how our two notions of stability relate to the one of Cohen et al. (2021).
They propose that the quantity which stabilizes around $2/\eta$ is, in our notation,
$\mathrm{IAS}\quad:=\quad\frac{\mathbb{E}_{\begin{subarray}{c}\text{sampling}\\ \text{batches}\end{subarray}}\left[\nabla f_{B}^{\top}\cdot\nabla^{2}f\cdot\nabla f_{B}\right]}{\|\nabla f\|^{2}}\quad=\quad\frac{\mathrm{tr}\left(\nabla^{2}f\cdot\mathbb{E}_{\begin{subarray}{c}\text{sampling}\\ \text{batches}\end{subarray}}\left[\nabla f_{B}\nabla f_{B}^{\top}\right]\right)}{\mathrm{tr}\left(\nabla f\nabla f^{\top}\right)}.$ (11)
The very natural way they come up with this quantity is by imposing the following inequality, similar to our Equation (10).
$\mathbb{E}_{\begin{subarray}{c}\text{sampling}\\ \text{batches}\end{subarray}}\left[\underbrace{f(x_{t+1})-f(x_{t})}\right]\ =\ -\eta\mathbb{E}_{\begin{subarray}{c}\text{sampling}\\ \text{batches}\end{subarray}}\left[\nabla\underbrace{f(x_{t})^{\top}}\nabla f_{B}(x_{t})\right]\,+\,\frac{\eta^{2}}{2}\mathbb{E}_{\begin{subarray}{c}\text{sampling}\\ \text{batches}\end{subarray}}\left[\nabla f_{B}(x_{t})^{\top}\cdot\underbrace{\nabla^{2}f(x_{t})}\nabla f_{B}(x_{t})\right]\ \leq\ 0.$ (12)
The difference between our Equation (10) and this equation above if that the terms in which we put a under brace are here evaluated on the whole dataset and not on the mini batch considered.
While the inequalities seem analogous and both equally motivated at first glance, we claim they carry an important difference. Enforcing Equation (12) does imply that in 1 step the loss does not go up. However, it does not imply that the loss will not go up eventually. The instability threshold discussed in Cohen et al. (2021), in the following work, and in our is stability about the whole trajectory, not about one precise step of the optimizer in consideration.

### v4

@@ A.3 Proof of Proposition 1
Expand $\|\nabla L_{B}\|^{2}=\|\nabla L+(\nabla L_{B}-\nabla L)\|^{2}$ and take $\mathbb{E}_{B}[\cdot]$. The cross term vanishes because $\mathbb{E}_{B}[\nabla L_{B}]=\nabla L$. ∎
Explicit variance.
Let $\Sigma(\theta):=\mathrm{Cov}_{(x,y)\sim\mathcal{D}}[\,\nabla L_{B}(\theta)\,]$, $B\sim\mathcal{P}_{b}$, $b=1$, be the covariance of a single sample gradient. Then, for i.i.d. batches of size $b$,
$\mathbb{E}_{B}\!\bigl{[}\|\nabla L_{B}-\nabla L\|^{2}\bigr{]}=\frac{1}{b}\,\operatorname{tr}\bigl{(}\Sigma(\theta)\bigr{)}.$
For a finite dataset of size $n$ sampled without replacement the factor $1/b$ is replaced by ${(n-b)}/\bigl{(}{b\,(n-1)}\bigr{)}$.
Stationary, small-step regime.
In Appendix A.3 we solved the discrete Lyapunov equation and obtained (first order in the step size $\eta$):
$\|\nabla L(\theta)\|^{2}=\frac{\eta}{b}\,\operatorname{tr}\!\bigl{(}\mathcal{H}\mathcal{K}^{\dagger}(\Sigma)\mathcal{H}\bigr{)}+\mathcal{O}(\eta^{2})$
where $\mathcal{K}(X)=\mathcal{H}X+X\mathcal{H}$ and $\mathcal{K}^{\dagger}$ is its Moore–Penrose inverse on $\operatorname{Im}\mathcal{H}\otimes\operatorname{Im}\mathcal{H}$. This implies that
$\frac{\|\nabla L\|^{2}}{\mathbb{E}_{B}[\|\nabla L_{B}\|^{2}]}\quad=\quad\eta\,\frac{\operatorname{tr}(\mathcal{H}\mathcal{K}^{\dagger}(\Sigma)\mathcal{H})}{\operatorname{tr}(\Sigma)+\eta\operatorname{tr}(\mathcal{H}\mathcal{K}^{\dagger}(\Sigma)\mathcal{H})}+\mathcal{O}(\eta^{2}).$ (9)
Spectral bounds on the bias–variance ratio.
Let the eigen–decomposition of the full-batch Hessian be $\mathcal{H}=\sum_{i=1}^{r}\lambda_{i}\,v_{i}v_{i}^{\!\top}$ with strictly positive eigen-values $0<\mu:=\lambda_{1}\leq\dots\leq\lambda_{r}=\lambda_{\max}$ and let $P_{+}:=\sum_{i=1}^{r}v_{i}v_{i}^{\!\top}$ be the projector onto $\operatorname{Im}\mathcal{H}$. Define the curved-subspace covariance $\Sigma_{+}:=P_{+}\Sigma P_{+}$. Then, up to the $\mathcal{O}(\eta^{2})$ term already displayed,
$\frac{\mu}{2}\,\operatorname{tr}(\Sigma_{+})\;\leq\;\operatorname{tr}\bigl{(}\mathcal{H}\,\mathcal{K}^{\dagger}(\Sigma)\,\mathcal{H}\bigr{)}\;\leq\;\frac{\lambda_{\max}}{2}\,\operatorname{tr}(\Sigma_{+})\;\leq\;\frac{\lambda_{\max}}{2}\,\operatorname{tr}(\Sigma).$ (10)
Consequently
Lemma 3 (Bound on the ratio).
In the oscillatory regime, we have
$\frac{\|\nabla L\|^{2}}{\mathbb{E}_{B}[\|\nabla L_{B}\|^{2}]}\;=\;\eta\,\frac{\operatorname{tr}(\mathcal{H}\mathcal{K}^{\dagger}(\Sigma)\mathcal{H})}{\operatorname{tr}(\Sigma)+\eta\operatorname{tr}(\mathcal{H}\mathcal{K}^{\dagger}(\Sigma)\mathcal{H})}+\mathcal{O}(\eta^{2})$
and
$\frac{\eta\,\mu}{2}\;\frac{\operatorname{tr}(\Sigma_{+})}{\operatorname{tr}(\Sigma)+\tfrac{\eta\lambda_{\max}}{2}\operatorname{tr}(\Sigma_{+})}\;\leq\;\frac{\|\nabla L\|^{2}}{\mathbb{E}_{B}[\|\nabla L_{B}\|^{2}]}\;\leq\;\frac{\eta\,\lambda_{\max}}{2}\;\frac{\operatorname{tr}(\Sigma_{+})}{\operatorname{tr}(\Sigma)+\tfrac{\eta\mu}{2}\operatorname{tr}(\Sigma_{+})}.$
If the gradient noise has no component in the null-space of $\mathcal{H}$ $(\Sigma_{+}=\Sigma)$,
$\frac{\eta\mu}{2+\eta\mu}\;\leq\;\varepsilon^{2}(\theta)\;\leq\;\frac{\eta\lambda_{\max}}{2+\eta\lambda_{\max}},\qquad\frac{\|\nabla L\|^{2}}{\mathbb{E}_{B}[\|\nabla L_{B}\|^{2}]})\in\bigl{[}\tfrac{\eta\mu}{2},\tfrac{\eta\lambda_{\max}}{2}\bigr{]}+\mathcal{O}(\eta^{2}).$
Proof.
Write $\Sigma=\sum_{i,j}s_{ij}\,v_{i}v_{j}^{\!\top}$. Because $\mathcal{K}^{\dagger}(v_{i}v_{j}^{\!\top})=(\lambda_{i}+\lambda_{j})^{-1}v_{i}v_{j}^{\!\top}$ for $\lambda_{i}+\lambda_{j}>0$,
$\mathcal{H}\,\mathcal{K}^{\dagger}(\Sigma)\,\mathcal{H}=\sum_{i,j}\frac{\lambda_{i}\lambda_{j}}{\lambda_{i}+\lambda_{j}}\,s_{ij}\;v_{i}v_{j}^{\!\top}.$
Taking the trace removes all off-diagonal terms and gives
$\operatorname{tr}\bigl{(}\mathcal{H}\,\mathcal{K}^{\dagger}(\Sigma)\,\mathcal{H}\bigr{)}=\sum_{i=1}^{r}\frac{\lambda_{i}^{2}}{2\lambda_{i}}\,s_{ii}=\frac{1}{2}\sum_{i=1}^{r}\lambda_{i}\,s_{ii}.$
Each coefficient $\lambda_{i}$ lies between $\mu$ and $\lambda_{\max}$, while $\sum_{i}s_{ii}=\operatorname{tr}(\Sigma_{+})$. This yields the two inequalities in (10). Insert them into the fraction $\eta\,\operatorname{tr}(\mathcal{H}\mathcal{K}^{\dagger}(\Sigma)\mathcal{H})/\bigl{[}\operatorname{tr}(\Sigma)+\eta\operatorname{tr}(\mathcal{H}\mathcal{K}^{\dagger}(\Sigma)\mathcal{H})\bigr{]}$ and simplify. ∎
A.5 Extending to Neural Networks
Differentiating Oscillations in Neural Network Optimization
Our analytical treatment of SGD on one-dimensional quadratic objectives leverages the simplicity of having a single curvature measure–the second derivative–which facilitates a precise landscape characterization and explicit stability conditions. However, extending this analysis to multidimensional quadratics already introduces significantly more intricate dynamics, necessitating advanced analytical frameworks as developed by [28, 29, 32]. Transitioning further to neural network optimization increases this complexity dramatically, since training predominantly occurs away from the manifold of minima, including the EoS-like instabilities themselves (as evidenced by the continuous reduction in loss)—and therefore requires to go beyond linear stability of quadratics near the manifold of minima.
Given the current absence of robust theoretical tools to comprehensively analyze such dynamics, distinguishing between curvature-driven and noise-driven oscillations necessitates empirical experimentation. Specifically, we probe the dynamics by systematically varying hyperparameters (e.g., step size or batch size), as illustrated in Figure 3, allowing us to differentiate curvature-induced (Type-2) oscillations from purely noise-induced (Type-1) oscillations (Figure 4).
Type-2 Oscillations Are Unique to NN Optimization
This complexity inherent in neural network optimization is not merely an analytical inconvenience; rather, it is intrinsically tied to the emergence and significance of Type-2 oscillations and EoS-style phenomena. Notably, Type-2 oscillations emerge naturally [^10] We define an as emerging naturally if it arises inherently from the training dynamics, and not a result of precisely-selected hyperparameters or initializations, reflecting a fundamental characteristic of the optimization process itself. Formally, it needs to happen over a range of hyperparameter choices and initializations. only in the case of neural network optimization, but not in the case of quadratic objectives. In the one-dimensional quadratic scenario analyzed previously, curvature-driven oscillations require the step size to precisely match the stability threshold $2/\lambda_{\max}$, or exceed it, in which case we have divergence—in either case, it means that optimization of quadratics does naturally enter a regime of instability. In contrast, neural network optimization uniquely exhibits progressive sharpening, a third-order derivative phenomenon [18], where curvature naturally increases during training. This progressive increase in curvature means that training with a fixed step size can transition into an EoS-like regime of instability without any explicit adjustment of the hyperparameters, and stay there due to self-stabilization effects [18]. Hence, Type-2 oscillations emerge naturally and robustly within neural network training dynamics due to this intrinsic change of the loss landscape. Consequently, Type-2 oscillations and EoS-like regimes are fundamentally driven by progressive sharpening, which doesn’t happen in quadratics, making it a purely neural network optimization phenomena.
Appendix B Quadratic Setting: Batch Sharpness and GNI
B.1 Setting: Data and risk.
Intuitively, we minimise a random quadratic loss whose curvature changes with each sample. Let $(H_{i},x_{i})_{i\geq 1}$ be i.i.d. with
$H_{i}\in\mathbb{R}^{d\times d},\;H_{i}\succeq 0,\quad x_{i}\in\mathbb{R}^{d},\quad\mathbb{E}\!\bigl{[}\lVert H_{i}\rVert_{\mathrm{F}}^{4}+\left\|x_{i}\right\|^{4}\bigr{]}<\infty.$
Define [^11] All eigenvalues of $H:=\mathbb{E}[H_{i}]$ are positive because $\lambda_{\min}(H)>0$ is assumed; the ‘$+$’ superscript in earlier drafts is therefore redundant.
$H:=\mathbb{E}[H_{i}],\quad G:=\mathbb{E}[H_{i}x_{i}],\quad\theta^{\star}:=H^{-1}G.$
The population risk is $L(\theta):=\mathbb{E}_{i}\bigl{[}(\theta-x_{i})^{\top}H_{i}(\theta-x_{i})\bigr{]}$.
B.2 SGD dynamics.
With constant stepsize $\eta$ satisfying
$0<\eta<\frac{2}{\lambda_{\max}(H)},$ (11)
the update
$\theta_{t+1}=\theta_{t}-2\eta\,H_{i_{t}}\bigl{(}\theta_{t}-x_{i_{t}}\bigr{)},$ (12)
where $i_{t}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}\mathrm{Unif}\{1,\dots,n\}$ and $i_{t}\!\perp\!\!\!\perp\!\theta_{t}$, generates a Markov chain that is geometrically ergodic [60]. Let $\pi_{\eta}$ denote its unique stationary law and write $m_{\eta}:=\mathbb{E}_{\pi_{\eta}}[\theta]$, $\Sigma_{\eta}:=\operatorname{Cov}_{\pi_{\eta}}(\theta)$.
Key intuition: the stochastic gradient $g(\theta,i):=2H_{i}(\theta-x_{i})$ is unbiased, $\mathbb{E}_{i}[g(\theta,i)]=\nabla L(\theta)=2H(\theta-\theta^{\star})$, so the mean iterate should coincide with $\theta^{\star}$.
B.3 Stationary mean and covariance.
The discussion in Appendix A implies that in this setting
Lemma 4 (Lyapunov solution).
For any stepsize satisfying (11),
$m_{\eta}$ $=\theta^{\star},$ (13)
$\Sigma_{\eta}$ $=\frac{\eta}{2}\,H^{-1}\Sigma_{g}H^{-1}+\mathcal{O}(\eta^{2}),\qquad\Sigma_{g}:=4\,\operatorname{Var}_{i}\!\bigl{[}H_{i}(\theta^{\star}-x_{i})\bigr{]},$ (14)
where the $\mathcal{O}(\eta^{2})$ constant depends polynomially on $\lambda_{\min}(H)^{-1}$ and on the $4^{\text{th}}$-moment bound above.
Sketch.
Set $\Delta_{t}:=\theta_{t}-\theta^{\star}$ and $\xi_{t}:=2H_{i_{t}}(\theta^{\star}-x_{i_{t}})$ (mean 00, variance $\Sigma_{g}$). Subtracting (12) at $\theta^{\star}$ gives
$\Delta_{t+1}=(I-2\eta H)\Delta_{t}-\eta\,\xi_{t}-2\eta\bigl{(}H_{i_{t}}-H\bigr{)}\Delta_{t}.$ (15)
Mean. Taking expectations and using $i_{t}\!\perp\!\!\!\perp\!\theta_{t}$, $\mathbb{E}[\Delta_{t+1}]=(I-2\eta H)\,\mathbb{E}[\Delta_{t}],$ whose unique fixed point is 00, proving $m_{\eta}=\theta^{\star}$ (13).
Covariance. Let $S_{t}:=\mathbb{E}[\Delta_{t}\Delta_{t}^{\top}]$. Multiplying (15) by its transpose and exploiting $\mathbb{E}[\Delta_{t}]=0$, one obtains the discrete Lyapunov recursion $S_{t+1}=(I-2\eta H)S_{t}(I-2\eta H)^{\top}+\eta^{2}\Sigma_{g}+R_{t}$ with $\|R_{t}\|=O(\eta^{3})$ thanks to $\mathbb{E}\|H_{i}\|_{\mathrm{F}}^{4}<\infty$. Passing to the limit and inverting $\bigl{[}I-(I-2\eta H)^{\otimes 2}\bigr{]}$ by a Neumann series yields (14). ∎
Take-away.
Constant-stepsize SGD is unbiased in the mean but has an $\mathcal{O}(\eta)$ stationary variance that blows up if $\eta$ is too large; the $2/\lambda_{\max}(H)$ is thus necessary for avoiding the blowup, although not sufficient.
B.4 Per-sample residuals.
Define, for any $\theta\in\mathbb{R}^{d}$,
$Y_{i}(\theta):=H_{i}(\theta-x_{i}),\quad\mu(\theta):=\mathbb{E}_{i}[Y_{i}(\theta)]=H\theta-G.$
At stationarity write $Y:=Y_{i}(\theta)$, $\mu:=\mu(\theta)$, and set
$\widetilde{Y}\;:=\;Y-\mathbb{E}_{i}[Y\mid\theta]\;=\;Y-\mu\quad\Longrightarrow\quad\mathbb{E}_{i}[\widetilde{Y}\mid\theta]=0.$
Finally define the scalars [^12] Notation: $\mathbb{E}_{i,\pi}$ integrates $i$ and $\theta\sim\pi_{\eta}$.
$\mathcal{A}$ $:=\mathbb{E}_{i,\pi}\bigl{[}Y^{\!\top}H_{i}Y_{i}\bigr{]},$$\quad\mathcal{B}$ $:=\mathbb{E}_{i,\pi}\bigl{[}Y^{\!\top}HY_{i}\bigr{]},$ $\quad\mathcal{C}$ $:=\mathbb{E}_{i,\pi}\bigl{[}\left\|Y_{i}\right\|^{2}\bigr{]},$ (16)
$C_{0}$ $:=\mathbb{E}_{i,\pi}\bigl{[}\left\|\widetilde{Y}_{i}\right\|^{2}\bigr{]},$ $\quad\mathcal{D}$ $:=\mu^{\top}\mathbb{E}_{i,\pi}\bigl{[}H_{i}Y_{i}\bigr{]}.$
Why these symbols? $\mathcal{A}$ is the average loss at stationarity; $\mathcal{B}$ and $\mathcal{C}$ act as mixed-moment controls that will upper-bound $\mathcal{A}$ via a trace inequality.
B.5 Bounding curvature–fluctuation interaction.
Lemma 5 (Cross term).
Let $\widetilde{\Delta}:=\mathbb{E}_{i,\pi}\bigl{[}\widetilde{Y}^{\!\top}(H_{i}-H)\widetilde{Y}\bigr{]}$ and $\Delta:=\mathcal{A}-\mathcal{B}=\mathbb{E}_{i,\pi}\bigl{[}Y^{\!\top}(H_{i}-H)Y\bigr{]}$. Then $\Delta:=\mathcal{A}-\mathcal{B}=\widetilde{\Delta}+2\mathcal{D}$,
$\widetilde{\Delta}\;\leq\;\sqrt{\mathbb{E}_{i}[\|H_{i}-H\|_{2}^{2}]}\;\sqrt{\mathbb{E}_{\pi}[\|\widetilde{Y}\|^{4}]},$ (17)
and
$\Delta\;\leq\;\sqrt{\mathbb{E}_{i}[\|H_{i}-H\|_{2}^{2}]}\;\sqrt{\mathbb{E}_{\pi}[\|\widetilde{Y}\|^{4}]+2C_{0}\lVert\mu\rVert^{2}+\lVert\mu\rVert^{4}}.$ (18)
Proof.
In the following two steps we establish the first inequality.
Step 1 (Cauchy - Schwarz). For any matrix $A$ and vector $v$, $|v^{\top}Av|\leq\|A\|_{2}\,\|v\|^{2}$. With $A:=H_{i}-H$ and $v:=Y$:
$|\,\widetilde{Y}^{\!\top}(H_{i}-H)\widetilde{Y}|\leq\|H_{i}-H\|_{2}\,\|\widetilde{Y}\|^{2}.$
Note that this is sharp if and only if $\widetilde{Y}$ aligns with the eigenvectors of the maximal eigenvalues of $H_{i}-H$.
Step 2 (Average and separate: Jensen). Independence of $i$ and $\theta$ at stationarity implies, taking the expectation that
$|\Delta|\leq\sqrt{\mathbb{E}_{i}[\|H_{i}-H\|_{2}^{2}]}\;\sqrt{\mathbb{E}_{\pi}[\|\widetilde{Y}\|^{4}]}.$
Second inequality. Note that